OpenStudy (pulsified333):

Let H(x)=F(G(x)) and J(x)=F(x)/G(x). Suppose F(2)=5, F′(2)=5 G(2)=9, G′(2)=0 G(−8)=2, G′(−8)=−6 then H(−8)= H′(−8)= J(2)= J′(2)=

OpenStudy (pulsified333):

I am thinking that product and quotient rule are both used but I am not sure

OpenStudy (pulsified333):

H(-8)=2 is wrong

OpenStudy (anonymous):

Why is it wrong?

OpenStudy (anonymous):

Oh My bad, it should be 5.

OpenStudy (anonymous):

Sorry about that.

OpenStudy (pulsified333):

its okay

Nnesha (nnesha):

for first one you need to apply chain rule there is function inside the function \[\huge\rm H(x)= F(\color{ReD}{G(x)})\] \[\rm =f'(g(x)) \cdot g'(x)\] i guess

Nnesha (nnesha):

plugin -8 for x to find h(-8)

OpenStudy (pulsified333):

@Nnesha which problem are you referring to?

Nnesha (nnesha):

to find H(-8) you need to use first function.

OpenStudy (pulsified333):

okay

OpenStudy (pulsified333):

For the method you mentioned I would have H(x)=(2*5)(-6)

OpenStudy (pulsified333):

@Nnesha

Nnesha (nnesha):

no that's not correct.

Nnesha (nnesha):

\[\huge\rm H(x)= F(\color{ReD}{G(x)})\] \[\rm =f'(\color{blue}{g(-8)}) \cdot g'(-8)\] G(-8)= ???

OpenStudy (pulsified333):

2

Nnesha (nnesha):

right so replace g(-8) by 2 \[\rm =f'(\color{blue}{2 }) \cdot g'(-8)\] make sense now ?

OpenStudy (pulsified333):

so its 5*-6

Nnesha (nnesha):

yep

OpenStudy (pulsified333):

this is for H(-8) correct?

Nnesha (nnesha):

yes

OpenStudy (pulsified333):

the answer is 5 but the answer I would of gotten would have been -30

Nnesha (nnesha):

ahh i see there are two different questions H(-8) AND H'(-8) the one we did is the 2nd one where we have to take derivative

OpenStudy (pulsified333):

oh okay

Nnesha (nnesha):

the first is simple: just pluggin n numbers

Nnesha (nnesha):

one*

OpenStudy (pulsified333):

I am confused on how to do the H'(-8) though

Nnesha (nnesha):

H'(-8) = means take the derivative when x=-8 so you have to apply the chain rule there

OpenStudy (pulsified333):

ok ill try it

OpenStudy (pulsified333):

how do i do the chain rule for this part? i know the derivative of x=-8 is -6, but im lost after that

OpenStudy (pulsified333):

@Nnesha

Nnesha (nnesha):

what do you mean "derivative of x = -8" ?? first you need to replace x by -8 \[\large\rm H(-8)= F(\color{blue}{G(-8)})\] and then take derivative

OpenStudy (kayders1997):

So when you do this you take the derivative of the outside leaving the inside the same like @Nnesha is saying so you will have F(G(x)) so take the derivative of the outside leaving the inside unchanged times the derivative of the anside and than stay with you unchanged term because that's the on,y part that has an x term for the first part of the equation

OpenStudy (kayders1997):

So what is the derivative of f and g

Nnesha (nnesha):

right side : there is a function inside the function so you have to do chain rule here is an example \[\large\rm f(\color{Red}{g(x)}) = f'(\color{red}{g(x)}) \cdot \color{blue}{g'(x)}\] take the derivative of outer function (inside function would stay the same ) and then multiply by the derivative of inner function

OpenStudy (pulsified333):

so does H'(-8)=-12?

Nnesha (nnesha):

how did you get that ??

OpenStudy (pulsified333):

oh wait. OMG! I am so dumb lol

OpenStudy (pulsified333):

H'(-8)=-30; I am sorry, I've had a long day :D

Nnesha (nnesha):

no you're not.

Nnesha (nnesha):

yep looks good o^_^o

OpenStudy (pulsified333):

Two college midterms in one day is exhausting

Nnesha (nnesha):

better than essay ;(